(2y ^ 2 + 3x) * d * x + 2xydy = 0 .
1
Expert's answer
2021-09-23T17:28:18-0400

"(2y^2+3x)dx+2xydy=0\\
M=2y^2+3xy,~~N=2xy\\
M_y=4y\neq N_x=2y\\
\text{Thus the differential equation is not exact.}\\
\text{But, }\\
\frac{M_y-N_x}{N}=\frac{1}{x} \text{ is a function of x. }\\
I.F= e^{\int\frac{1}{x}dx}=e^{\ln x}=x\\
\text{Multiply the original equation by the integrating factor}\\
(2xy^2+3x^2)dx+2x^2ydy=0\\
M_y=4xy=N_x=4xy\\
\text{Then, it is exact}\\
F_x=M\\
F_x=2xy^²+3x^2\\
\text{Integrate both sides with respect to } x\\
F=x^2y^2+x^3+\Phi(y)\\
\text{Differentiate w.r.t } y\\
F_y=2x^2y+\Phi'(y)\\
2x^2y=2x^2y+\Phi'(y)\\
\Phi'(y)=0\\
\Phi(y)=C\\
\implies F=x^2y^2+x^3+C"